OpenStudy (anonymous):

Use the Substitution method to solve the system of equations. x + y = 20 30x + 25y = 550

OpenStudy (anonymous):

First turn one of the equations into slope-intercept form by (I would suggest the first one) and then plug in what you get for y into the other equation and solve for x

OpenStudy (anonymous):

x + y = 20 y=-x+20 30x + 25(-x+20)= 550 30x-25x+500=550 5x+500=550

OpenStudy (anonymous):

thank you.

OpenStudy (anonymous):

anytime :)